OpenStudy (anonymous):

y = -x + x2 can some one plz tell me how to graph this?

OpenStudy (anonymous):

y=-x+x^2 rewriting it. y= x^2 - x can you see the highest power/exponent is 2? This means that it will be a parabola. and the coefficient(number) in front of x^2 term which is positive 1, tells us it is a positive parabola. 1x^2 = x^2 there is actually a 1 still in front of the x^2 but we just don't show it. Find the x-intercepts,y-intercepts and turning point(vertex) and join them like a curve. Note: there are not always x-intercepts or y-intercepts it depends.

OpenStudy (anonymous):

 That is a basic positive parabola, of the original form x^2 note that the coefficient is positive one, hence that the graph goes upwards. If it is negative, it will go downwards.